OpenStudy (anonymous):

Help! Find the zeros of f(x)=x^3-7x^2-30x.

OpenStudy (lgbasallote):

to find the zeros you need to factor out and equate f(x) to 0 \[x^3 - 7x^2 - 30x = 0\] \[x(x^2 - 7x - 30) = 0\] \[x(x-10)(x+3) = 0\] now solve for x \[x =0; \; x = 10 ; \; x= -3\] does that help?